how to write equation forline passing through (-1,20) perpendicular to graph 4x+20y=15
Answer by mananth(16949) 
 4 x + 20 y = 15
Find the slope of this line
20 y = -4 x + 15
Divide by 20
y = - 1/ 5 x + 3/4
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 1/5
The slope of a line perpendicular to the above line will be the negative reciprocal 5
m1*m2=-1
The slope of the required line will be 5
m= 5 ,point ( -1 , 20 )
Find b by plugging the values of m & the point in
y=mx+b
20 = -5 + b
b= 25
m= 5
The required equation isy = 5 x+25
m.ananth@hotmail.ca
|
|
|
| |